The following gives information about the proportion of a sample that agree with a certain statement. Use StatKey or other technology to find a confidence interval at the given confidence level for the proportion of the population to agree, using percentiles from a bootstrap distribution. StatKey tip: Use "CI for Single Proportion" and then "Edit Data" to enter the sample information.

Find a confidence interval if 35 agree in a random sample of 100 people.

Click here to access StatKey.

Round your answers to two decimal places.

The confidence interval is                to
Answer #1

Ans- Below are given the required answer.

Confidence Interval For Proportion
CI = p ± Z a/2 Sqrt(p*(1-p)/n)))
x = Mean
n = Sample Size
a = 1 - (Confidence Level/100)
Za/2 = Z-table value
CI = Confidence Interval
Mean(x)=35
Sample Size(n)=100
Sample proportion = x/n =0.35
Confidence Interval = [ 0.35 ±Z a/2 ( Sqrt ( 0.35*0.65) /100)]
= [ 0.35 - 1.96* Sqrt(0.00228) , 0.35 + 1.96* Sqrt(0.00228) ]
= [ 0.25651,0.44349]